OpenStudy (anonymous):

Adult tickets cost $18 and children cost $12. A total of 1460 tickets were sold. $22,752 was collected from the sale. How many adult tickets were sold? How do you solve this? I have gotten this far: Let x = adult tickets Let y= children tickets 18x + 12y = 22,752 x+y=1460 Did I set this up right and what next?

OpenStudy (anonymous):

You did set this up right. Next, isolate both y's and use the formula y1 = y2

OpenStudy (anonymous):

You did this right**

OpenStudy (anonymous):

Do you still need help?

OpenStudy (anonymous):

Yes I thought I had it but I am still confused

OpenStudy (anonymous):

So both of your formulas are good

OpenStudy (anonymous):

I just don't know what to do next.

OpenStudy (anonymous):

I have trouble when there are 2 variables

OpenStudy (anonymous):

Now let's get y1 = y2 so 12y = -18x + 22 752 y = -1.5x +1896 y= -x +1460 -x + 1460 = -1.5x +1896

OpenStudy (anonymous):

From here, you can find the value of x adn with the value of x, you can find the value of y.

OpenStudy (anonymous):

awesome

OpenStudy (anonymous):

x=872 whoo hoo!